A class has 35 children there are 5 more girls than boys. How many are boys and how many are girls
Found 3 solutions by rfer, greenestamps, amalm06:
Answer by rfer(16322)

x=girls
x+5=boys
-----------------
x+x+5=35
2x=35-5
x=30/2
x=15
x+5=20

Answer by greenestamps(13215)


Often, the algebra we do to solve a problem like this is exactly what we could do to solve the problem using logical analysis.

Compare the solution below using logical analysis to the very good algebraic solution provided by the other tutor:

(1) Have 5 of the girls step out of the room.
(2) That leaves equal numbers of boys and girls, with a total of 30 students.
(3) So there are 15 boys and 15 girls.
(4) Bring the 5 girls back into the room.
(5) There are now 15 boys and 20 girls.

Do you see that this is EXACTLY what the formal algebra does?

I am not saying that you should use logical analysis instead of formal algebra. Quite the opposite -- I am saying that you should understand WHY the solution using formal algebra makes sense.

Answer by amalm06(224)

Let x=# of boys, y=#of girls
x+y=35
x=y-5
So y-5+y=35
2y=40
y=20
20 girls, 15 boys